# How Do Social Bots Participate in Misinformation Spread? A Comprehensive Dataset and Analysis

### v2

@@ Abstract
The social media platform is an ideal medium to spread misinformation, where social bots might accelerate the spread. This paper is the first to explore the interplay between social bots and misinformation on the Sina Weibo platform. We construct a large-scale dataset that contains annotations of misinformation and social bots. From the misinformation perspective, this dataset is multimodal, containing 11,393 pieces of misinformation and 16,416 pieces of real information. From the social bot perspective, this dataset contains 65,749 social bots and 345,886 genuine accounts, where we propose a weak-supervised annotator to annotate automatically. Extensive experiments prove that the dataset is the most comprehensive, misinformation and real information are distinguishable, and social bots have high annotation quality. Further analysis illustrates that: (i) social bots are deeply involved in information spread; (ii) misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots amplify this phenomenon; and (iii) social bots generate similar content aiming to manipulate public opinions.
@@ A Comprehensive Dataset and Analysis
Herun Wan1        Minnan Luo1        Zihan Ma1 Guang Dai2        Xiang Zhao3 1Xi’an Jiaotong University    2SGIT AI    3National University of Defense Technology wanherun@stu.xjtu.edu.cn
@@ 1 Introduction
Social media platforms, like $\mathbb{X}$ (Twitter) and Weibo, have become major information sources, and information spreads faster than traditional media. Due to the nature of such platforms, there have been attempts to disseminate misinformation, which could polarize society (Azzimonti and Fernandes, 2023) and impact the economy (Zhou et al., 2024). Meanwhile, besides attracting genuine users, the straightforward operation of social media also makes it an ideal breeding ground for fully automatic programs. Such programs, also known as social bots, could help publish various news and provide emergency help. While social bots can be considered a mere technological instrument, their malicious misuse could seriously compromise online safety. They are proven behind many online perils, including election interference (Ng et al., 2022) and hate speech propaganda (Stella et al., 2019). Social bots are natural message amplifiers (Caldarelli et al., 2020), increasing the risk of spreading misinformation (Huang et al., 2022). Namely, misinformation and social bots are two major factors endangering online security, and they would work together to amplify their impact as Figure 1 shows.
@@
Figure 1: An example of social bots amplifying information spread. Social bots would publish similar content to amplify misinformation spread and manipulate public sentiment and stance, increasing harmfulness.
@@
Researchers make efforts to fight the never-ending plague of misinformation and malicious social bots. They mainly propose automatic detectors to identify misinformation (Shu et al., 2019) and social bots (Yang et al., 2022). Meanwhile, researchers also explore how different types of content (Nan et al., 2021) or propagation patterns (Vosoughi et al., 2018) influence misinformation spread. From the social bot perspective, bot communities (Tan et al., 2023b) and bots’ repost behaviors (Elmas et al., 2022) have been investigated. While existing works have provided valuable insights into investigating misinformation and social bots, relatively little attention has been paid to the interplay between them.
@@
In this paper, we aim to bridge the gap of existing works, investigating the interplay between misinformation and social bots. We construct a novel dataset named MisBot [^1] We have published MisBot in GitHub., which simultaneously contains information and annotations of misinformation and social bots (§2). Specifically, we first define the structure of MisBot. We then collect misinformation from Weibo’s official management center [^2] https://service.account.weibo.com/, being available for users who have logged in.. After that, we propose a simple but effective debiasing method to mitigate entity biases and collect real information from two credible sources. We finally employ a weak-supervised annotator to annotate users who participate in the information spread automatically. From the misinformation perspective, MisBot contains multiple modalities, including post content, comments, repost messages, images, and videos. MisBot includes 11,393 pieces of misinformation and 16,416 pieces of real information. From the social bot perspective, MisBot includes 952,955 users participating in the information spread, covering 65,749 annotated social bots and 345,886 genuine accounts. Extensive experiments (§3) prove that (i) MisBot is the most comprehensive and the only one with misinformation and social bot annotations, (ii) misinformation and real information are distinguishable, where a simply detector achieves 95.2% accuracy, and (iii) MisBot has high social bot annotation quality, where human evaluations prove it. Further analysis illustrates (§4) that (i) social bots are deeply involved in information spread, where 29.3% users who repost misinformation are social bots; (ii) misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots amplify this phenomenon; and (iii) social bots generate similar content aiming to manipulate public opinions, including sentiments and stances.
@@ 2 MisBot Dataset
The collection process of MisBot consists of four components: (i) Data Structure, which defines the dataset structure; (ii) Misinformation Collection, which collects multiple modalities in misinformation; (iii) Real Information Collection, which collects real information without entity biases; and (iv) Weak-Supervised User Annotation, which trains a weak-supervised model to automatically annotate related accounts.
@@ 2.1 Data Structure
Unlike news platforms, users publish posts on Weibo, thus, we annotate user posts as misinformation or real information, where the main language is Chinese. From the online information perspective, each instance is represented as $A=\{s,\mathcal{G}_{\textit{repost}},\mathcal{G}_{\textit{comment}},I,V,y\}$. It contains textual content $s$, repost graph $\mathcal{G}_{\textit{repost}}$, comment graphs $\mathcal{G}_{\textit{comment}}$, images $I$, videos $V$, and corresponding label $y$. From the account perspective, each instance is represented as $U=\{F,T,y\}$. It contains the attribute set $F$, published posts $T$, and the corresponding label $y$. Meanwhile, if a user reposts, comments, or likes a post, we consider it to be participating in the spread of information. We provide example instances in MisBot in Appendix A.1.
@@ 2.2 Misinformation Collection
We collect posts flagged as misinformation from Weibo’s official management center [^3] We present the platform overview in Appendix A.2.. This platform presents posts containing misinformation judged by platform moderators or police. Besides, it provides a brief judgment to explain why the post is flagged as misinformation, which provides a basis for identifying topics of misinformation. We present an example in Appendix A.3. We have collected all the misinformation since this platform was established. Specifically, the misinformation collected was published between April 2018 and April 2024. We spent about 10 months collecting 11,393 pieces of misinformation.
@@ 2.3 Real Information Collection
Existing misinformation datasets generally suffer from potential data bias (Chen et al., 2023), especially entity biases (Zhu et al., 2022), which would influence models’ generalization ability to unseen data. Thus, we design an entity debiasing method to mitigate entity biases. We first employ a keyphrase extractor (Liang et al., 2021) to obtain key entities from each misinformation. After filtering [^4] We present the filter strategy in Appendix A.4 ., we obtain 1,961 entities. We finally query the key entities in trusted sources to get real information. To ensure the authenticity and diversity of real information, we collect real information from two sources:
  • Verified news media is an official news account certified by the Weibo platform, which contains a red “verified” symbol and a verified reason [^5] We provide verified news accounts in Appendix A.5..
@@
We discuss the credibility of these two sources and provide examples and detailed query methods in Appendices A.6 and A.7. We obtained 8,317 and 8,099 pieces of real information, respectively.
2.4 Weak-Supervised User Annotation
Manual annotation or crowd-sourcing is labor-intensive and not feasible with large-scale datasets. Meanwhile, to ensure the scalability of MisBot, we propose a weak supervision learning strategy, enabling the automatic annotation. The construction of the weak-supervised annotator contains three phases: (i) preprocessing phase; (ii) training phase; and (iii) inference phase.
@@ 2.4 Weak-Supervised User Annotation
This phase preprocesses the training dataset for the weak-supervised annotator. We first collected 100,000 random accounts. Due to the randomness, these accounts could represent the entire Weibo environment, ensuring the diversity of accounts. We employ crowd-sourcing to annotate them, where the human annotators are familiar with social media. We summarize the general criteria for identifying a social bot on Weibo and write a guideline document for human annotators based on the criteria. We provide the criteria and guideline document in Appendix A.8. To ensure the annotation quality, each human annotator is required to read the document before annotating. Each annotator should also annotate 20 standard accounts that are obviously social bots or genuine accounts. Annotators who achieve more than 80% accuracy on standard accounts are reliable. We ensure that each account is annotated by three reliable human annotators. We provide the number of human annotators, the salary of annotators, and the annotation cost in Appendix A.9. We employ major voting to obtain the final annotations in this phase.
We filter in active accounts [^6] We provide the filter strategy based on human annotators’ feedback in Appendix A.10. and focus on them in MisBot, where there are three main reasons:
@@
  • Content-based detectors leverage NLP technologies to analyze user-generated content, where we employ name, description, and posts. We employ encoder-based language models, including BERT (Devlin et al., 2019) and DeBERTa (He et al., 2021) to identify social bots.
  • Ensemble detectors combine user attributes and generated content to identify social bots.
We provide the descriptions and settings of experts in Appendix A.11. We create an 8:1:1 split for the users from the preprocessing phase as train, validation, and test sets to train each expert.
@@
This phase annotates an unseen account based on the predictions from multiple experts. To ensure annotation quality, we filter in the experts achieving 80% accuracy, which is the same standard as human annotators, on the validation set. After that, a conventional method to integrate multiple predictions is to employ majority voting or train an MLP classifier on the validation set (Bach et al., 2017; Feng et al., 2022). Since the likelihood from classifiers may not accurately reflect true probabilities (Guo et al., 2017), also known as miscalibrated, we calibrate the likelihoods before the ensemble. We employ temperature scaling and select the best temperature on the validation set [^7] We provide the temperature settings in Appendix A.12.. We finally average the calibrated likelihoods to obtain the final annotations. Among the 952,955 accounts that participate in information spread in MisBot, 411,635 are active, of which 65,749 are social bots and 345,886 are genuine accounts.
@@ 3 Basic Analysis of MisBot
MisBot is the most comprehensive
@@
Misinformation and real information in MisBot are distinguishable
@@
From the data distribution perspective, we first explore the differences in multi-modal content consistency between misinformation and real information. We employ three metrics [^8] We provide the calculation formula in Appendix B.1.: (i) Text, where the higher value means the text is more consistent with the global texts; (ii) Similarity, where the higher value means the text and image are more consistent; and (iii) Image, where the higher value means the image is more consistent with the global images. Figure 2 visualizes the joint distribution of these metrics. For Text, real information illustrates higher textual content consistency, where the majority lies at 0.6, while misinformation shows lower, where some even show negative values. We speculate that real information would follow a certain pattern, while misinformation would not. For Similarity, real information illustrates relatively more consistency than misinformation, where real information might publish text-related images and misinformation might not. Thus, we could conclude that misinformation and real information present different multi-modal content consistency, being distinguishable. To further capture the image differences between misinformation and real information, we present the distribution of image categories and sentiments in Figure 14 in Appendix B.2. It illustrates that misinformation and real information present distinct distributions [^9] We provide the experiment settings and detailed analysis in Appendix B.2..
@@
w/o Interaction $\mathop{81.6^{\star}_{\pm 4.5}}\limits_{14.2\%\downarrow}$ $\mathop{77.3^{\star}_{\pm 4.1}}\limits_{16.2\%\downarrow}$ $\mathop{64.4^{\star}_{\pm 5.9}}\limits_{31.3\%\downarrow}$ $\mathop{97.3^{\star}_{\pm 1.2}}\limits_{7.0\%\uparrow}$
w/o Vision $\mathop{94.1^{\star}_{\pm 0.5}}\limits_{1.1\%\downarrow}$ $\mathop{90.3^{\star}_{\pm 1.0}}\limits_{2.2\%\downarrow}$ $\mathop{94.2^{\dagger}_{\pm 1.2}}\limits_{0.4\%\uparrow}$ $\mathop{86.8^{\star}_{\pm 2.1}}\limits_{4.6\%\downarrow}$
w/o Extra $\mathop{78.5^{\star}_{\pm 5.2}}\limits_{17.5\%\downarrow}$ $\mathop{74.5^{\star}_{\pm 4.3}}\limits_{19.3\%\downarrow}$ $\mathop{60.6^{\star}_{\pm 6.0}}\limits_{35.4\%\downarrow}$ $\mathop{97.3^{\star}_{\pm 1.0}}\limits_{6.9\%\uparrow}$
@@
From the misinformation detector perspective, we design a simple misinformation detector to verify whether the detector could identify misinformation in MisBot [^10] We provide the details of this model in Appendix B.3.. We present the performance of the detector and variants in Table 2. This simple detector achieves remarkable performance, where the f1-score reaches 92.3%. The ideal performance proves that misinformation and real information are easily distinguished by a machine detector, which helps explore the differences between social bots in spreading misinformation and real information. Meanwhile, the detector without interaction drops to 77.3% on f1-score, illustrating the effectiveness of user reactions, which coincides with our speculation that social bots might have different social patterns in misinformation and real information [^11] We provide more analysis about the ablation study in Appendix B.4..
MisBot has high social bot annotation quality, where the weak-supervised annotator is reliable
The construction of the weak-supervised annotator contains three phases, where the results from each phase are reliable:
@@
  • Training phase. We employed multiple detectors aiming to identify various social bots. To ensure the weak-supervised annotator’s credibility, we filtered in detectors achieving 80% accuracy and obtained 4 detectors [^12] We provide the performance of each detector and the corresponding temperature in Appendix B.5.. The accuracy on the test set reaches 85.03%, which is higher than TwiBot-20 (Feng et al., 2021b) and TwiBot-22 (Feng et al., 2022), illustrating credibility.
  • Inference phase. To evaluate annotation quality in MisBot, we randomly sample 50 social bots and 50 genuine accounts and manually annotate them through a human expert. The Cohen’s Kappa between the human expert and automatic annotator is 0.74, showing good agreement.

### v3

@@ Abstract
Social media platforms provide an ideal environment to spread misinformation, where social bots can accelerate the spread. This paper explores the interplay between social bots and misinformation on the Sina Weibo platform. We construct a large-scale dataset that includes annotations for both misinformation and social bots. From the misinformation perspective, the dataset is multimodal, containing 11,393 pieces of misinformation and 16,416 pieces of verified information. From the social bot perspective, this dataset contains 65,749 social bots and 345,886 genuine accounts, annotated using a weakly supervised annotator. Extensive experiments demonstrate the comprehensiveness of the dataset, the clear distinction between misinformation and real information, and the high quality of social bot annotations. Further analysis illustrates that: (i) social bots are deeply involved in information spread; (ii) misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots would amplify this phenomenon; and (iii) social bots generate similar content aiming to manipulate public opinions.
@@ A Comprehensive Dataset and Analysis
Herun Wan1, 2 Minnan Luo1, 2, * Zihan Ma1, 3 Guang Dai4 Xiang Zhao5 1School of Computer Science and Technology, Xi’an Jiaotong University, China 2Ministry of Education Key Laboratory of Intelligent Networks and Network Security, China 3Shaanxi Province Key Laboratory of Big Data Knowledge Engineering, China 4SGIT AI Lab, State Grid Corporation of China 5National University of Defense Technology wanherun@stu.xjtu.edu.cn minnluo@xjtu.edu.cn https://whr000001.github.io/MisBot/
**footnotetext: Corresponding author
@@ 1 Introduction
Social media platforms, like $\mathbb{X}$ (Twitter) and Weibo, have become major information sources, and information spreads faster than traditional media. Due to the nature of such platforms, there have been attempts to disseminate misinformation, which could polarize society (Azzimonti and Fernandes, 2023) and impact the economy (Zhou et al., 2024). Meanwhile, besides attracting genuine users, the social platform also becomes an ideal breeding ground for malicious social bots (Cresci, 2020) due to the straightforward operation. Social bots are proven behind many online perils, including election interference (Ng et al., 2022) and hate speech propaganda (Stella et al., 2019). Social bots are natural message amplifiers (Caldarelli et al., 2020), increasing the risk of spreading misinformation (Huang et al., 2022). Namely, misinformation and social bots are two major factors harming online security. They might work together to amplify negative impact, where Figure 1 presents an example.
@@
Figure 1: An example of social bots participating in information spread. Social bots would publish similar content to manipulate public sentiment and stance, leading to a shift in public opinion.
@@
Researchers make efforts to fight the never-ending plague of misinformation and malicious social bots. They mainly propose automatic detectors to identify misinformation (Shu et al., 2019) and social bots (Yang et al., 2022). Meanwhile, researchers also explore how different types of content (Nan et al., 2021) or propagation patterns (Vosoughi et al., 2018) influence misinformation spread. From the social bot perspective, bot communities (Tan et al., 2023b) and bots’ repost behaviors (Elmas et al., 2022) have been investigated. While many works have provided valuable insights into investigating misinformation and social bots, relatively little attention (Wang et al., 2018; Himelein-Wachowiak et al., 2021) has been paid to the interplay between them.
@@
This paper aims to bridge the gap of existing works, exploring the interplay between misinformation and social bots. We propose MisBot [^1] The main language of MisBot is Chinese., a dataset which simultaneously contains information and annotations of misinformation and social bots (§2). Specifically, we first define the structure of MisBot. We then collect misinformation from Weibo’s official management center [^2] https://service.account.weibo.com/, being available for users who have logged in.. After that, we collect real information from two credible sources to ensure the authenticity. We finally propose a weakly supervised annotator to label the users involved in the dissemination of information. From the misinformation perspective, MisBot contains multiple modalities, including post content, comments, repost messages, images, and videos. MisBot includes 11,393 misinformation instances and 16,416 real information instances. From the social bot perspective, MisBot includes 952,955 users participating in the information spread, covering 65,749 annotated social bots and 345,886 genuine accounts. Extensive experiments (§3) prove that (i) MisBot is the most comprehensive and the only one with misinformation and social bot annotations, (ii) misinformation and real information are distinguishable, where a simple detector achieves 95.2% accuracy, and (iii) MisBot has high social bot annotation quality, where human evaluations prove it. Further analysis illustrates (§4) that (i) social bots are deeply involved in information spread, where 29.3% users who repost misinformation are social bots; (ii) misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots amplify this phenomenon; and (iii) social bots generate similar content aiming to manipulate public opinions, including sentiments and stances.
@@ 2 MisBot Dataset
The collection process of MisBot consists of four components: (i) Data Structure defines the dataset structure; (ii) Misinformation Collection collects multiple modalities in misinformation; (iii) Real Information Collection collects real information from two sources; and (iv) Weakly Supervised User Annotation trains a weakly supervised annotator to automatically annotate accounts.
@@ 2.1 Data Structure
Users publish posts to spread information on the Weibo platform, thus, we annotate user posts as misinformation or real information. From this perspective, each instance is represented as $A=\{s,\mathcal{G}_{\textit{repost}},\mathcal{G}_{\textit{comment}},I,V,y\}$. It contains textual content $s$, repost graph $\mathcal{G}_{\textit{repost}}$, comment graphs $\mathcal{G}_{\textit{comment}}$, images $I$, videos $V$, and corresponding label $y$. From the account perspective, each instance is represented as $U=\{F,T,y\}$. It contains the attribute set $F$, published posts $T$, and the corresponding label $y$. Meanwhile, we believe a user participates in the spread of a post if this user reposts, comments, or likes this post. Some cases in MisBot are provided in Appendix A.1.
@@ 2.2 Misinformation Collection
We collect posts flagged as misinformation from Weibo’s official management center, where we provide the platform overview in Appendix A.2 for readers who cannot log in. This platform presents posts containing misinformation judged by platform moderators or police. Besides, it provides a brief judgment to explain why the post is flagged as misinformation, which provides a basis for identifying topics of misinformation. An example is provided in Appendix A.3. We have collected all the misinformation since this platform was established. Specifically, the misinformation collected was published between April 2018 and April 2024. We spent about 10 months collecting 11,393 pieces of misinformation.
@@ 2.3 Real Information Collection
Existing misinformation datasets generally suffer from potential data bias (Chen et al., 2023), especially entity biases (Zhu et al., 2022). It means that the entity distributions in misinformation and real information differ, influencing models’ generalization ability to unseen data. Thus, we design an entity debiasing method to mitigate entity biases. We first employ a keyphrase extractor (Liang et al., 2021) to obtain key entities from each misinformation. After filtering uncommon entities, we get 1,961 entities, where we present the filter rules in Appendix A.4. We finally query the key entities using the Weibo search engine in trusted sources to get real information. An overview of the search engine is provided in Appendix A.5 for readers who cannot log in. To ensure the authenticity and diversity of real information, we collect real information from two sources:
  • Verified news media is an official news account certified by the Weibo platform, which contains a red “verified” symbol and a verified reason, where we provide the statistics of the verified news accounts in Appendix A.6.
@@
Due to the moderation of Weibo, we assume these two sources are truthful, where we discuss it further in Appendix A.7 and quantitatively prove it in §3. We obtained 8,317 and 8,099 pieces of real information, respectively.
2.4 Weakly Supervised User Annotation
Manual annotation or crowd-sourcing is labor-intensive and not feasible with large-scale datasets. Meanwhile, to ensure the scalability of MisBot, we propose a weakly supervised learning strategy, enabling automatic annotation. The construction of the weakly supervised annotator contains (i) preprocessing, (ii) training, and (iii) inference phases.
@@ 2.4 Weakly Supervised User Annotation
This phase aims to collect the training dataset for the weakly supervised annotator. We first collected 100,000 random accounts. Due to the randomness, these accounts could represent the entire Weibo environment, ensuring the diversity of accounts. We employ crowd-sourcing to annotate them, where the human annotators are familiar with social media. Following existing works (Feng et al., 2021b, 2022), we summarize a brief criteria for identifying a social bot on Weibo and write a guideline document for human annotators, where we provide the document in Appendix A.8. Notably, social bot annotation is subjective, where the average Fleiss’ Kappa is 0.4281 as shown in §3. Thus, we do not directly define what a social bot is, but only provide a brief guideline document and cases. Inspired by existing work (Feng et al., 2021b), we determine 20 standard accounts that are easy to identify. Each annotator should also annotate 20 standard accounts, and annotators who achieve more than 80% accuracy on standard accounts are reliable. We ensure that each account is annotated by three reliable human annotators. We totally recruited 315 annotators and spent 60,000 yuan and 60 days, where we provide details in Appendix A.9. We employ major voting to obtain the final annotations in this phase.
Based on human annotators’ feedback, we filter in active accounts in MisBot, where we provide the filter rules in Appendix A.10. We focus on active accounts for three reasons:
@@
  • Content-based detectors encode user-generated textual content, where we employ name, description, and posts. We employ encoder-based language models, including BERT (Devlin et al., 2019) and DeBERTa (He et al., 2021) to obtain textual representations and employ an MLP layer to identify social bots.
  • Ensemble detectors concatenate the attribute and textual representations and employ an MLP layer to identify social bots.
The descriptions and settings of experts are provided in Appendix A.11. We create an 8:1:1 split for the users from the preprocessing phase as train, validation, and test sets to train each expert.
@@
This phase annotates accounts based on the predictions from multiple experts. To ensure annotation quality, we filter in the experts achieving 80% accuracy, which is the same standard as human annotators, on the validation set. After that, a conventional method to integrate multiple predictions is to employ majority voting or train an MLP classifier on the validation set (Bach et al., 2017; Feng et al., 2022). Since the likelihood from classifiers may not accurately reflect true probabilities (Guo et al., 2017), also known as miscalibrated, we calibrate the likelihoods before the ensemble. We employ temperature scaling (Guo et al., 2017) and select the best temperature on the validation set, where we provide the temperature settings in Appendix A.12. We finally average the calibrated likelihoods to obtain the final annotations. Among the 952,955 accounts that participate in information spread in MisBot, 411,635 are active, of which 65,749 are social bots and 345,886 are genuine accounts.
@@ 3 Basic Analysis of MisBot
MisBot is the most comprehensive.
@@
Misinformation and real information in MisBot are distinguishable.
@@
From the data distribution perspective, we first explore the differences in content consistency between misinformation and real information. We employ three metrics: (i) Text to evaluate the text consistency of a specific instance and all instances; (ii) Image to evaluate the image consistency of a specific instance and all instances; and (iii) Similarity to evaluate the consistency of text and image in a specific instance. We provide the calculation formula in Appendix B.1 and present the joint distributions in Figure 2. It illustrates that misinformation and real information present distinct consistency. Specifically, real information presents higher Text and Similarity. Namely, we could conclude that misinformation and real information are distinguishable in terms of consistency.
To further capture the image differences between misinformation and real information, we present the distribution of image categories and sentiments in Figure 15 in Appendix B.2. It illustrates that misinformation and real information present distinct distributions. Specifically, real information would contain more neutral images while misinformation would contain more screenshots.
@@
[…]
@@
From the misinformation detector perspective, we design a simple misinformation detector to verify whether the detector could identify misinformation in MisBot, where we provide the details of this model in Appendix B.3. We present the performance of the detector and ablation variants in Table 2. This simple detector achieves remarkable performance, where the accuracy reaches 95.2%. The ideal performance proves that misinformation and real information are easily distinguished by a machine detector, which helps explore the differences between social bots in spreading misinformation and real information. Meanwhile, the detector without interaction drops to 77.3% on f1-score, illustrating the effectiveness of user reactions, which coincides with our speculation that social bots might have different social patterns in misinformation and real information. We also provide a complete analysis of the ablation study in Appendix B.4.
MisBot has high social bot annotation quality, where the weakly supervised annotator is reliable.
The construction of the weakly supervised annotator contains three phases, where we have proven that each phase is reliable:
@@
  • Training phase. We employed multiple detectors aiming to identify various social bots. To ensure the annotator’s credibility, we filtered in detectors achieving 80% accuracy and obtained 4 detectors. The accuracy on the test set reaches 85.03%, which is higher than TwiBot-20 (Feng et al., 2021b) and TwiBot-22 (Feng et al., 2022), illustrating credibility. We also provide the performance of each detector and the corresponding temperature in Appendix B.5.
  • Inference phase. We randomly sample 50 social bots and 50 genuine accounts in MisBot and manually annotate them through a human expert. The Cohen’s Kappa between the human expert and the automatic annotator is 0.74, showing good agreement.